OpenStudy (anonymous):

Given the equation y − 3 = 1/2(x + 6) in point-slope form, identify the equation of the same line in slope-intercept form.

OpenStudy (anonymous):

Slope intercept of a equation of a line is y=mx+c

OpenStudy (anonymous):

So, try to take y − 3 = 1/2(x + 6) in the form of y=mx+c